Morning Litho status report 6/27

Gary Sosa gsosa at stanford.edu
Mon Jun 27 07:22:39 PDT 2011


Good morning All....

    Here is the morning Litho status report.

Nanospec 2(Litho)- Wavelength counter not working. Will check out on 
morning rounds.

Karl Suss 1- Scheduled downtime today. 6 month PM to be performed by vendor.


*Note* *1*: There will be an ASML Stepper maintenance training class 
from 6/13 through 7/1. During certain parts
            of the day the stepper will be unavailable as the trainers 
will be using the tool for the "hands-on" portion
            of the training class. There will be a large group gathered 
around the stepper during this time and the area
            around the stepper will be blocked off to normal traffic. 
Please observe the barricades during this  time
            period. We apologize for the inconvenience and thank you for 
your patience and understanding.


          Please check back in Coral for the latest tool updates and 
current equipment status.

                     Thanks...  The Litho Team


*/Photolithography Area Status/*

	

	

*/Monday, June 27, 2011/*

	

	

	

	

*/General Information / Lab Activities/*

	

	

	

ASML Stepper maintenance training class from 6/13 through 7/1

*/Exposure Tools/**//*

	

Status

	

Comments

	

Priority

ASML Stepper

	

*PROBLEM*

	

Intermittant carrier inspection

	

2

EVAlign

	

*PROBLEM*

	

Non-criticals only

	

3

EVAlign2

	

*UP*

	

	

EVG-Imprint

	

*UP*

	

	

Karl Suss 1

	

*UP*

	

6 Month PM today

	

Karl Suss 2

	

*UP*

	

	

*/Manual Coaters/*

	

Status

	

	

EVGSpraycoat

	

*UP*

	

	

Headway

	

*UP*

	

	

Laurell

	

*UP*

	

	

*/Coaters/*

	

Process Status

	

	

SVGCoat

	

3612 Process

	

*Up*

	

	

220-3 Process

	

*Up*

	

	

220-7 Process

	

*Up*

	

	

SVGCoat2

	

3612 Process

	

*Up*

	

	

955-0.7 Process

	

*Up*

	

	

3617 Process

	

*Up*

	

	

*/Developers/*

	

Status

	

	

SVGDEV

	

*UP*

	

	

SVGDEV2

	

*UP*

	

	

*/Ovens / Hotplates/*

	

Status

	

	

BLUEM

	

*UP*

	

	

Blue M 90 C.

	

*UP*

	

	

Blue M 110 C

	

*UP*

	

	

Blue M 150 C

	

*UP*

	

	

Fusion

	

*UP*

	

	

UV-Cure

	

*UP*

	

	

White-Oven

	

*UP*

	

	

YES Vapor Prime

	

*UP*

	

	

*/Hotplates- /*

	

	

	

Laurel-left

	

*UP*

	

	

Laurel- right

	

*UP*

	

	

Headway

	

*UP*

	

	

Torrey Pines- left

	

*UP*

	

	

90 Deg C

	

*UP*

	

	

115 Deg C

	

*UP*

	

	

Torrey Pines- Right

	

*UP*

	

	

*/Metrology/Inspection/*

	

Status

	

	

Alpha-Step 500

	

*UP*

	

	

Nanospec 2-Litho

	

*DOWN*

	

Wavelength counter problem

	

1

Nanospec 1-Etch

	

*UP*

	

	

P2 Profiler

	

*UP*

	

	

*/Microscpoes-/*

	

	

	

#1 Leica/Analytical

	

*UP*

	

	

#2 Ergolux/Drytek2

	

*UP*

	

	

#3 Olympus/Litho

	

*UP*

	

	

#4 Nikon Optiphot/Litho

	

*UP*

	

	

#5 Nikon/Litho

	

*UP*

	

	

#6 Leitz/Bosch

	

*UP*

	

	

#7 Leitz/Diffusion

	

*UP*

	

	

#8 Olympus/Litho

	

*UP*

	

	

#9 Olympus/Litho

	

*UP*

	

	

*/Miscellaneous/*

	

Status

	

	

EVBond

	

*UP*

	

	

KSBonder

	

*UP*

	

	

KSCoat

	

*UP*

	

	

Litho SRD

	

*UP*

	

	

Mask Cleaner

	

*UP*

	

	

	

	

	


-------------- next part --------------
An HTML attachment was scrubbed...
URL: <http://snf.stanford.edu/pipermail/litho/attachments/20110627/d6597612/attachment.html>


More information about the litho mailing list